We want to design an airport runway with the following specifications. The lowest acceleration of a plane should be  




4m/s2




and its take-off speed is 75 m/s. How long would the runway be to allow the planes to accelerate through it?





1
Expert's answer
2022-06-06T08:53:33-0400
"v=v_0+at"

"t=\frac{v-v_0}{a}=\frac{75-0}{4}=19\:\rm s"